### v1

@@ 1.1 Sample
In a previous paper Chen et al. 2023 (hereafter as Chen2023; [21]), we selected giant planets around single Sun-like stars in the Galactic disk from the planet host catalog of Planet Across Space and Time (PAST) I [26] and divide the planetary sample into three subsamples, i.e., radial velocity (RV), ground-based transit (GT) and space-based transit (ST). Then we constructed parent samples for the three subsamples (see § 3 of supporting information (SI) of Chen2023 [21]). Specifically, the RV parent sample were constructed by using data from Keck, Lick, HAPRS [48, 49, 50]. The ST parent sample were constructed by using data from LAMOST-Gaia-Kepler catalog [27]. For the GT subsample, we conducted their parent sample primary based on the information from (Super)WASP survey [8, 51], which searched stars with V-band magnitudes in the range of $\sim 8-13$ [8]. For homogeneity, we only adopted Tycho-2 catalog, which is only 99% complete to V-band magnitude of $V\lesssim 11$ [52].
In this work, to cover the entire magnitude range of SuperWASP and further expand the GT planetary subsample, we utilize the Gaia DR2 catalog, which provides parallax, celestial coordinates (RA and DEC), proper motions, and magnitudes. We then select stars from the Gaia DR2 catalog using similar criteria as SuperWASP [51]:
@@
(2) Dwarf stars by excluding giant stars with $\log g<4.0$. Note this criterion is not exactly same as the reduced proper motion (RPM) method but it offers a more direct approach with the same physical meaning. We then collect RV data from APOGEE, RAVE, LAMOST and Gaia with the same criteria as PAST I. Using the astrometric and RV data, we calculate their kinematic properties (i.e., Galactic position, velocity, and component). We make the quality control and keep Sun-like stars in the Galactic disk with the same criteria as described in § 1 of the SI of Chen2023 [21]. The final parent GT stellar sample contains 915,925 stars. We then crossmatch the parent stellar sample with planets discovered by transit method and ground-based facilities to obtain the planetary sample. To ensure uniformity in the RV follow-up of SuperWASP transit candidates, we retain only hot Jupiters with high signal-to-red noise ratios (see § 3.2.1 of SI of Chen2023 [21] for more details). The number of hot Jupiters in the GT sample increases from 43 (in Chen 2023) to 92.
In sum, based on the kinematic catalogs of PAST series [26, 27], we select giant planets around single Sun-like stars and construct their parent stellar samples. The final sample consists of:
@@ 1.2 Deriving the $F_{\rm HJ}$ by combining RV, GT and ST data
    We calculate the detection efficiencies and geometric effects for stars and planets for the three subsamples with the same procedure described in § 3.1-3.3 of SI of Chen2023 [21]. We then obtain the $N_{\rm p}$} and $N_{\rm eff}$} of the ten bins for the three subsamples.
@@
    We eliminate the effect of stellar $\rm[Fe/H]$ by unifying to solar metallicities (i.e., $\rm[Fe/H]=0$). It has been found that the frequencies of giant planets are positively correlated with stellar $\rm[Fe/H]$ as $F_{\rm HJ}\propto 10^{\beta\times\rm[Fe/H]}$ [21]. Thus the stellar $\rm[Fe/H]$ induce a difference in $F_{\rm HJ}$ by a factor of:
@@ 1.3.1 $F_{\rm HJ}$ is generally declining with age
We divide the whole planetary sample into ten bins according to the relative membership probability for the thick disk to thin disk $TD/D$. Then we calculate the frequencies of hot Jupiters by adopting the methods described in § 1.2. For each bin, the kinematic age $t$ is calculated from the vertical velocity dispersion using the refined age-velocity dispersion relation (AVR) of PAST I [26], which gives
@@
where $t$ is stellar age, $\sigma$ is the velocity dispersion. $k$ and $b$ are two coefficients for AVR (listed in Table 5 of [26]). The uncertainties of kinematic ages $t$ are calculated by means of error propagation considering the uncertainties of vertical velocity dispersion and the AVR coefficients.
@@
For each model, we fit the relation between $F_{\rm HJ}$ and $\bar{t}$ (Gyr) with the Levenberg-Marquardt algorithm (LMA). In order to avoid being trapped in some local best fits, we randomly generate 1,000 sets of initial starting guesses of the fitting parameters before LMA fitting. We then select the one with lowest residual sum of squares (RSS) from the 1,000 local best-fits as the global best fit. In order to compare the global best fits of the two models, we calculate the Akaike information criterion (AIC) [53] for each model. The resulting AIC values are 93.2-93.2 and 113.5-113.5 for the constant and exponential model, respectively. Therefore a constant model can be confidently ruled out compared with the exponential model with AIC score difference $\rm\Delta AIC>10$.
@@
which is consistent with the results derived from the three subsamples in Chen2023 [21].
@@ 1.3.3 The two populations of hot Jupiters of different origin channels of different timescales
$t_{\rm in}$ and $t_{\rm cir}$ are estimated using the following equations [54, 13]:
@@
where $M_{*}$ and $M_{\rm p}$ are the masses of the host star and the planet. $R_{*}$ and $R_{\rm p}$ are the radii of the host star and the planet. $a$ and $P$ are the orbital semi-major axis and period of the planet. The subscripts $\odot$ and $\rm J$ denote the Sun and Jupiter. $Q_{\rm p}$ is the planetary tidal quality factor and we here adopt a fiducial value often used for close Jovian planets as $10^{6}$ referring to [55, 56]. $Q^{{}^{\prime}}_{*}$ is the stellar tidal quality factor and we adopt the best fits from the observational results by comparing with the theoretical predictions (see later analyses in § 3.1).
In Extended Data Fig. 2, we also plot the cumulative distribution of the difference between the kinematic age and tidal evolution timescale, $t-t_{\rm tide}$ for the ‘late-arrived’ population. As can been, the majority of ‘late-arrived’ hot Jupiters should take $\gtrsim 1$ Gyr to reach their current orbits since the formation of their host stars, which provides a strong constraint to the ‘late’ model. As discussed later, we adopt the secular chaos model from [31] as the ‘late’ model. Compared to the arrived timescale via secular chaos model [31], $t-t_{\rm tide}$ from the observational sample seems to be larger. This is expected since our selection criteria for the ‘late-arrived’ observational sample are rather strict. As we mentioned before, a portion of ‘late-arrived’ hot Jupiters (e.g., $t<t_{\rm tide}$) may not be identified. Thus, we generated a synthetic population of 10,000 hot Jupiters with the arrival timescale distribution as [31]. We then calculate their inspiral timescales. Subsequently, we assigned ages and uncertainties to these synthetic hot Jupiters by randomly sampling from the observed ‘late-arrived’ hot Jupiter population. Then, we selected survived synthetic hot Jupiters that can be classified as ‘late-arrived’. Finally, we conducted a two-sample KS test between the arrival timescales of synthetic and observed hot Jupiters that can be selected as ‘late-arrived’, yielding a p-value of 0.3499. The above consistence further supports the secular chaos as the ‘Late’ model.
@@ 2.1 The origin of hot Jupiters
At present, there are mainly three classes of origin theories for the hot Jupiters: in-situ formation, disk migration, and high-eccentricity migration [2]. In the in-situ formation scenario, hot Jupiters form at their current locations through gravitational instability [57] and/or core accretion [5, 58]. It was believed that in-situ formation is difficult to operate at close-in orbit [59, 60], but some recent studies suggest that it may be feasible under some certain conditions [3, 4], which however could not explain the vast majority of hot Jupiter population. Thus some form of migrations must have occurred for most of hot Jupiters after their initial formation. To date, two primary migration mechanisms have been proposed: disk migration and high-eccentricity migration. In the disk migration scenario, hot Jupiters form several astronomical units (AU) away and migrate inward to $\lesssim 0.1$ AU under torques from the gaseous protoplanetary disk [5, 61, 6]. For the high-eccentricity migration scenario, hot Jupiters are driven into high-eccentricity orbits under close encounters between planets [7], Kozai–Lidov interactions [62, 63] induced by a distant binary companion star/massive planet on an inclined orbit [28, 9], secular chaos in multi-planetary systems in mildly inclined and eccentric orbits [10, 31] or planet-planet coplanar secular [11] and then moved from a large separation to several hundredths of AU [64, 65].
@@
Hot Jupiters of different origins are expected to have distinct formation timescales. For in-situ formation and gas disk migration, hot Jupiters must form or migrate to their current locations before the dissipation of the gas disk, i.e., $\lesssim 1-10$ Myr [66]. In contrast, for the high-eccentricity migration, the timescale varies significantly for different mechanisms. Specifically, in the case of planet-planet scattering, the eccentricities of hot Jupiters increase via a random walk process, with their periapsides shrinking to $\sim 0.1$ AU over a timescale of thousands of years [67]. For hot Jupiters formed via secular Kozai–Lidov interactions by inclined massive planets [9], most of them are tidally ensnared on their first close approach to the host star with a timescale ranging from $10^{4}-10^{8}$ years. Afterward, they are tidally circularized to orbits of $\lesssim 0.1$ AU within $\sim 1$–100 Myr (summarized in Fig. 3 and § 2.3.1.3 of [2]). Whereas, under the secular chaos in multi-planetary systems, the migration timescale of hot Jupiter can span the entire lifetime of the host star, i.e., $\sim 0.1-10$ Gyr [10, 31].
@@
Alternatively, the planet-star kozai effect can also deliver hot Jupiters under certain conditions [28, 29]. The planet eccentricity and periapse oscillate with a typical timescale as [68]:
@@
where $m_{1}$, $m_{2}$, $m_{3}$ are the masses of central stars, planets and outer companions, respectively. $P_{\rm in}$ and $P_{\rm out}$ are the orbital periods of the giant planet and companion. $e_{\rm out}$ is the eccentricity of the companion. We then exclude planets that are likely dynamically unstable according to the following formula: $\frac{a_{\rm in}}{a_{\rm out}}>0.330-0.417e_{\rm out}+0.069e_{\rm out}^{2}$ [69]. Furthermore, because Kozai cycles are driven by weak gravitational forces from the outer binary, they can be easily suppressed by general relativity pericenter precession in the inner binary [29]. Therefore, the Kozai effects can only occur when their timescale is shorter than the pericenter precession timescale (following Equation (23) in [29]). To obtain the distributions of the Kozai timescales, we adopt the following conditions: the distributions of orbital period, eccentricity and mass ratio of outer companions referring fig. 13, 15 and 16 of [70], and the distribution of orbital semi-major axes of giant planets between 1 and 5 au based on the power law distribution from [71]. Then we calculate with Equation S13 and keep those shorter than the corresponding pericenter precession timescale. Extended Data Fig. 1 shows the distribution of the planet-star Kozai timescale. As shown, the Kozai timescales are mainly between $10^{4}-10^{8}$ years, which is comparable to the results of previous studies [69]. The majority (94.28%) of the Kozai timescales are less than 0.1 Gyr, with nearly all (99.71%) being less than 1 Gyr.
@@
  • ‘Early’ model, i.e., in-situ foramtion, disk migration and high-eccentricity migration via planet-planet scattering [7]/Kozai-Lidov interactions [9]/coplanar secular [11]. Their formation timescales, typically within several tenths of Gyr, are comparable to (or even shorter than) the typical uncertainties in kinematic age estimates and are significantly shorter than the evolutionary timescales of our sample (i.e., $\sim 0.5$–10 Gyr). Therefore, given the current limitations in age accuracy, we collectively group the aforementioned mechanisms into a single category.
  • ‘Late’ model, i.e., High-eccentricity migration via secular chaos in multi-planetary systems, which could form hot Jupiters over a relatively long timescale extending to several Gyrs. Previous studies show that the stellar Kozai effect can also deliver hot Jupiters over Gyr under certain conditions [29, 10], but here we do not consider it as the main channel for the following reason: (1) the fraction of hot Jupiter host stars with binary companions capable of inducing Kozai–Lidov oscillations is at most $16\%\pm 5\%$ (when overcoming the general relativity pericenter precession) and will further decrease to $6\%\pm 2\%$ assume an isotropic distribution of the inclinations of stellar companions [30]. (2) The stellar kozai would mainly deliver hot Jupiters with timescales $\lesssim$ tenths of Gyr. (3) the potential binaries have been excluded in our sample.

### v2

@@ 1.1 Sample
In a previous paper Chen et al. 2023 (hereafter as Chen2023; [2023PNAS..12004179C]), we selected giant planets around single Sun-like stars in the Galactic disk from the planet host catalog of Planet Across Space and Time (PAST) I [2021ApJ...909..115C] and divide the planetary sample into three subsamples, i.e., radial velocity (RV), ground-based transit (GT) and space-based transit (ST). Then we constructed parent samples for the three subsamples (see § 3 of supporting information (SI) of Chen2023 [2023PNAS..12004179C]). Specifically, the RV parent sample were constructed by using data from Keck, Lick, HAPRS [2014ApJS..210....5F, 2017AJ....153..208B, 2020A&A...636A..74T]. The ST parent sample were constructed by using data from LAMOST-Gaia-Kepler catalog [2021AJ....162..100C]. For the GT subsample, we conducted their parent sample primary based on the information from (Super)WASP survey [2003ApJ...585.1056G, 2007ASPC..366..187W], which searched stars with V-band magnitudes in the range of $\sim 8-13$ [2003ApJ...585.1056G]. For homogeneity, we only adopted Tycho-2 catalog, which is only 99% complete to V-band magnitude of $V\lesssim 11$ [2000A&A...355L..27H].
In this work, to cover the entire magnitude range of SuperWASP and further expand the GT planetary subsample, we utilize the Gaia DR2 catalog, which provides parallax, celestial coordinates (RA and DEC), proper motions, and magnitudes. We then select stars from the Gaia DR2 catalog using similar criteria as SuperWASP [2007ASPC..366..187W]:
@@
(2) Dwarf stars by excluding giant stars with $\log g<4.0$. Note this criterion is not exactly same as the reduced proper motion (RPM) method but it offers a more direct approach with the same physical meaning. We then collect RV data from APOGEE, RAVE, LAMOST and Gaia with the same criteria as PAST I. Using the astrometric and RV data, we calculate their kinematic properties (i.e., Galactic position, velocity, and component). We make the quality control and keep Sun-like stars in the Galactic disk with the same criteria as described in § 1 of the SI of Chen2023 [2023PNAS..12004179C]. The final parent GT stellar sample contains 915,925 stars. We then crossmatch the parent stellar sample with planets discovered by transit method and ground-based facilities to obtain the planetary sample. To ensure uniformity in the RV follow-up of SuperWASP transit candidates, we retain only hot Jupiters with high signal-to-red noise ratios (see § 3.2.1 of SI of Chen2023 [2023PNAS..12004179C] for more details). The number of hot Jupiters in the GT sample increases from 43 (in Chen 2023) to 92.
In sum, based on the kinematic catalogs of PAST series [2021ApJ...909..115C, 2021AJ....162..100C], we select giant planets around single Sun-like stars and construct their parent stellar samples. The final sample consists of:
@@ 1.2 Deriving the $F_{\rm HJ}$ by combining RV, GT and ST data
    We calculate the detection efficiencies and geometric effects for stars and planets for the three subsamples with the same procedure described in § 3.1-3.3 of SI of Chen2023 [2023PNAS..12004179C]. We then obtain the $N_{\rm p}$} and $N_{\rm eff}$} of the ten bins for the three subsamples.
@@
    We eliminate the effect of stellar $\rm[Fe/H]$ by unifying to solar metallicities (i.e., $\rm[Fe/H]=0$). It has been found that the frequencies of giant planets are positively correlated with stellar $\rm[Fe/H]$ as $F_{\rm HJ}\propto 10^{\beta\times\rm[Fe/H]}$ [2023PNAS..12004179C]. Thus the stellar $\rm[Fe/H]$ induce a difference in $F_{\rm HJ}$ by a factor of:
@@ 1.3.1 $F_{\rm HJ}$ is generally declining with age
We divide the whole planetary sample into ten bins according to the relative membership probability for the thick disk to thin disk $TD/D$. Then we calculate the frequencies of hot Jupiters by adopting the methods described in § 1.2. For each bin, the kinematic age $t$ is calculated from the vertical velocity dispersion using the refined age-velocity dispersion relation (AVR) of PAST I [2021ApJ...909..115C], which gives
@@
where $t$ is stellar age, $\sigma$ is the velocity dispersion. $k$ and $b$ are two coefficients for AVR (listed in Table 5 of [2021ApJ...909..115C]). The uncertainties of kinematic ages $t$ are calculated by means of error propagation considering the uncertainties of vertical velocity dispersion and the AVR coefficients.
@@
For each model, we fit the relation between $F_{\rm HJ}$ and $\bar{t}$ (Gyr) with the Levenberg-Marquardt algorithm (LMA). In order to avoid being trapped in some local best fits, we randomly generate 1,000 sets of initial starting guesses of the fitting parameters before LMA fitting. We then select the one with lowest residual sum of squares (RSS) from the 1,000 local best-fits as the global best fit. In order to compare the global best fits of the two models, we calculate the Akaike information criterion (AIC) [CAVANAUGH1997201] for each model. The resulting AIC values are 93.2-93.2 and 113.5-113.5 for the constant and exponential model, respectively. Therefore a constant model can be confidently ruled out compared with the exponential model with AIC score difference $\rm\Delta AIC>10$.
@@
which is consistent with the results derived from the three subsamples in Chen2023 [2023PNAS..12004179C].
@@ 1.3.3 The two populations of hot Jupiters of different origin channels of different timescales
$t_{\rm in}$ and $t_{\rm cir}$ are estimated using the following equations [2009MNRAS.395.2268B, 2012MNRAS.423..486L]:
@@
where $M_{*}$ and $M_{\rm p}$ are the masses of the host star and the planet. $R_{*}$ and $R_{\rm p}$ are the radii of the host star and the planet. $a$ and $P$ are the orbital semi-major axis and period of the planet. The subscripts $\odot$ and $\rm J$ denote the Sun and Jupiter. $Q_{\rm p}$ is the planetary tidal quality factor and we here adopt a fiducial value often used for close Jovian planets as $10^{6}$ referring to [2004ApJ...610..477O, 2014ApJ...787...27Q]. $Q^{{}^{\prime}}_{*}$ is the stellar tidal quality factor and we adopt the best fits from the observational results by comparing with the theoretical predictions (see later analyses in § 3.1).
In Extended Data Fig. 2, we also plot the cumulative distribution of the difference between the kinematic age and tidal evolution timescale, $t-t_{\rm tide}$ for the ‘late-arrived’ population. As can been, the majority of ‘late-arrived’ hot Jupiters should take $\gtrsim 1$ Gyr to reach their current orbits since the formation of their host stars, which provides a strong constraint to the ‘late’ model. As discussed later, we adopt the secular chaos model from [2017MNRAS.464..688H] as the ‘late’ model. Compared to the arrived timescale via secular chaos model [2017MNRAS.464..688H], $t-t_{\rm tide}$ from the observational sample seems to be larger. This is expected since our selection criteria for the ‘late-arrived’ observational sample are rather strict. As we mentioned before, a portion of ‘late-arrived’ hot Jupiters (e.g., $t<t_{\rm tide}$) may not be identified. Thus, we generated a synthetic population of 10,000 hot Jupiters with the arrival timescale distribution as [2017MNRAS.464..688H]. We then calculate their inspiral timescales. Subsequently, we assigned ages and uncertainties to these synthetic hot Jupiters by randomly sampling from the observed ‘late-arrived’ hot Jupiter population. Then, we selected survived synthetic hot Jupiters that can be classified as ‘late-arrived’. Finally, we conducted a two-sample KS test between the arrival timescales of synthetic and observed hot Jupiters that can be selected as ‘late-arrived’, yielding a p-value of 0.3499. The above consistence further supports the secular chaos as the ‘Late’ model.
@@ 2.1 The origin of hot Jupiters
At present, there are mainly three classes of origin theories for the hot Jupiters: in-situ formation, disk migration, and high-eccentricity migration [2018ARA&A..56..175D]. In the in-situ formation scenario, hot Jupiters form at their current locations through gravitational instability [2007prpl.conf..607D] and/or core accretion [1996Natur.380..606L, 2014prpl.conf..619C]. It was believed that in-situ formation is difficult to operate at close-in orbit [2005ApJ...621L..69R, 2006ApJ...648.. [^6] R], but some recent studies suggest that it may be feasible under some certain conditions [2016ApJ...829..114B, 2016ApJ...817L..17B], which however could not explain the vast majority of hot Jupiter population. Thus some form of migrations must have occurred for most of hot Jupiters after their initial formation. To date, two primary migration mechanisms have been proposed: disk migration and high-eccentricity migration. In the disk migration scenario, hot Jupiters form several astronomical units (AU) away and migrate inward to $\lesssim 0.1$ AU under torques from the gaseous protoplanetary disk [1996Natur.380..606L, 2008ApJ...685..584I, 2014prpl.conf..667B]. For the high-eccentricity migration scenario, hot Jupiters are driven into high-eccentricity orbits under close encounters between planets [1996Sci...274..954R], Kozai–Lidov interactions [1962AJ.....67..591K, 1962P&SS....9..719L] induced by a distant binary companion star/massive planet on an inclined orbit [2003ApJ...589..605W, 2011Natur.473..187N], secular chaos in multi-planetary systems in mildly inclined and eccentric orbits [2011ApJ...735..109W, 2017MNRAS.464..688H] or planet-planet coplanar secular [2015ApJ...805...75P] and then moved from a large separation to several hundredths of AU [2014A&A...562A..71B, 2019A&A...631A..47K].
@@
Hot Jupiters of different origins are expected to have distinct formation timescales. For in-situ formation and gas disk migration, hot Jupiters must form or migrate to their current locations before the dissipation of the gas disk, i.e., $\lesssim 1-10$ Myr [1986ApJ...309..846L]. In contrast, for the high-eccentricity migration, the timescale varies significantly for different mechanisms. Specifically, in the case of planet-planet scattering, the eccentricities of hot Jupiters increase via a random walk process, with their periapsides shrinking to $\sim 0.1$ AU over a timescale of thousands of years [2012ApJ...751..119B]. For hot Jupiters formed via secular Kozai–Lidov interactions by inclined massive planets [2011Natur.473..187N], most of them are tidally ensnared on their first close approach to the host star with a timescale ranging from $10^{4}-10^{8}$ years. Afterward, they are tidally circularized to orbits of $\lesssim 0.1$ AU within $\sim 1$–100 Myr (summarized in Fig. 3 and § 2.3.1.3 of [2018ARA&A..56..175D]). Whereas, under the secular chaos in multi-planetary systems, the migration timescale of hot Jupiter can span the entire lifetime of the host star, i.e., $\sim 0.1-10$ Gyr [2011ApJ...735..109W, 2017MNRAS.464..688H].
@@
Alternatively, the planet-star kozai effect can also deliver hot Jupiters under certain conditions [2003ApJ...589..605W, 2007ApJ...669.1298F]. The planet eccentricity and periapse oscillate with a typical timescale as [1998MNRAS.300..292K]:
@@
where $m_{1}$, $m_{2}$, $m_{3}$ are the masses of central stars, planets and outer companions, respectively. $P_{\rm in}$ and $P_{\rm out}$ are the orbital periods of the giant planet and companion. $e_{\rm out}$ is the eccentricity of the companion. We then exclude planets that are likely dynamically unstable according to the following formula: $\frac{a_{\rm in}}{a_{\rm out}}>0.330-0.417e_{\rm out}+0.069e_{\rm out}^{2}$ [2007ApJ...670..820W]. Furthermore, because Kozai cycles are driven by weak gravitational forces from the outer binary, they can be easily suppressed by general relativity pericenter precession in the inner binary [2007ApJ...669.1298F]. Therefore, the Kozai effects can only occur when their timescale is shorter than the pericenter precession timescale (following Equation (23) in [2007ApJ...669.1298F]). To obtain the distributions of the Kozai timescales, we adopt the following conditions: the distributions of orbital period, eccentricity and mass ratio of outer companions referring fig. 13, 15 and 16 of [2010ApJS..190....1R], and the distribution of orbital semi-major axes of giant planets between 1 and 5 au based on the power law distribution from [2008PASP..120..531C]. Then we calculate with Equation S13 and keep those shorter than the corresponding pericenter precession timescale. Extended Data Fig. 1 shows the distribution of the planet-star Kozai timescale. As shown, the Kozai timescales are mainly between $10^{4}-10^{8}$ years, which is comparable to the results of previous studies [2007ApJ...670..820W]. The majority (94.28%) of the Kozai timescales are less than 0.1 Gyr, with nearly all (99.71%) being less than 1 Gyr.
@@
  • ‘Early’ model, i.e., in-situ foramtion, disk migration and high-eccentricity migration via planet-planet scattering [1996Sci...274..954R]/Kozai-Lidov interactions [2011Natur.473..187N]/coplanar secular [2015ApJ...805...75P]. Their formation timescales, typically within several tenths of Gyr, are comparable to (or even shorter than) the typical uncertainties in kinematic age estimates and are significantly shorter than the evolutionary timescales of our sample (i.e., $\sim 0.5$–10 Gyr). Therefore, given the current limitations in age accuracy, we collectively group the aforementioned mechanisms into a single category.
  • ‘Late’ model, i.e., High-eccentricity migration via secular chaos in multi-planetary systems, which could form hot Jupiters over a relatively long timescale extending to several Gyrs. Previous studies show that the stellar Kozai effect can also deliver hot Jupiters over Gyr under certain conditions [2007ApJ...669.1298F, 2011ApJ...735..109W], but here we do not consider it as the main channel for the following reason: (1) the fraction of hot Jupiter host stars with binary companions capable of inducing Kozai–Lidov oscillations is at most $16\%\pm 5\%$ (when overcoming the general relativity pericenter precession) and will further decrease to $6\%\pm 2\%$ assume an isotropic distribution of the inclinations of stellar companions [2016ApJ...827....8N]. (2) The stellar kozai would mainly deliver hot Jupiters with timescales $\lesssim$ tenths of Gyr. (3) the potential binaries have been excluded in our sample.